OpenStudy (anonymous):

my watch showed the correct time at 12 noon, but then the battery started to run down untill it eventually stopped completely. between 12 noon and when it stopped, it lost 20 minutes per hour on average.it now shows 6 pm but it stopped 4 hours ago. what is the correct time now? a.1am b.2am c.12am d.11am

OpenStudy (amistre64):

its 903pm in florida ;)

OpenStudy (anonymous):

i know that its 9 03pm there coz i m from there too

OpenStudy (anonymous):

but i need answer for the above question can anyone help me plz?

OpenStudy (amistre64):

my watch showed the correct time at 12 noon, but then the battery started to run down untill it eventually stopped completely. between 12 noon and when it stopped, it lost 20 minutes per hour on average. it now shows 6 pm but it stopped 4 hours ago. 12 + (1 -1/3) + (1 -1/3) + (1 -1/3+... 12 + 6 hours is 18 hours total 12 + (1 -1/3)x = 18

OpenStudy (amistre64):

x - (1/3)x = 6 (2/3)x = 6 x = 9 it would take 9 hours for it to reach 6 o'clock

OpenStudy (amistre64):

12 + 9 + 4 = 15 hours that have passed since noon

OpenStudy (anonymous):

the correct answeer is option no a .1 am

OpenStudy (anonymous):

but couldnt solve ..

OpenStudy (amistre64):

yeah, 1 am. 12 + 9 = 9pm + 4 hours = 1am

OpenStudy (amistre64):

do you see how I did it?

OpenStudy (anonymous):

how 1/3 cam there.

OpenStudy (amistre64):

20 minutes is 1/3 of an hour right?

OpenStudy (anonymous):

ye

OpenStudy (amistre64):

12 + 1 would be a normal hour correct? but it loses 1/3 hour each time. 12 + (1 -1/3) is how the watch runs...

OpenStudy (amistre64):

it ran like that till it hit 6 oclock; which is 18 hours in total for the day right?

OpenStudy (amistre64):

12 + (1 -1/3)(x hours) = 18 (1 -1/3)(x hours) = 18-12 = 6 (2/3)(x hours) = 6 hours x = 6(3)/2 = 18/2 = 9 9 hours it ran slow then died. 12 + 9 = 9pm + 4hours = 1am

OpenStudy (anonymous):

oh i got direclty howz that like.......its 3 hrs lost and it is added to 6pm and it becomes 9pm and it stoppped 4 hrs ago so we can add 9pm+4hrs=1 am.....thanks buddy am i right??